OpenStudy (anonymous):

Just curious: How did 0 over 0 becomes undefined? If we suppose that it has an equivalent value and when we multiply both sides by 0 it will be 0=0. so why undefined?

OpenStudy (amistre64):

by the mere fact that its not able to be defined a defininte value

OpenStudy (amistre64):

if 0/0 = some umber k 0/0 = k would suggest that there is some unique value k such that 0 = 0k

OpenStudy (kainui):

The reason 0/0 is undefined is because you can interpret it as many things. Calculus basically looks at this.

OpenStudy (anonymous):

hmmmm. so zero isn't considered as a unique value?

OpenStudy (amistre64):

of course not... 0*1 = 0 0*2 = 0 0*234958723498673894689035672372306 = 0 0*pi = 0 0*e = 0 etc ...

OpenStudy (amistre64):

do all those values of 0*k suggest that they are all equal to 0?

OpenStudy (anonymous):

okay. thanks. :) @amistre64

OpenStudy (anonymous):

Anything multiplied by 0 is 0. Therefore, 0/0 is undefined cuz it can have many solutions.

OpenStudy (amistre64):

undefined does not mean useless :) it just means that there is no way to define a numerical value to it

OpenStudy (kainui):

Similarly we can also think about 1 to any positive power. 1^0, 1^1, 1^2, etc...

OpenStudy (anonymous):

aaaaahhh. now I get it. thanks sir @amistre64 and others that replied. thank you. :D

OpenStudy (kainui):

Sorry, more like 1 to any integer power, not positive.

OpenStudy (kainui):

The connection is log(1)=0 for every base, that's really all I'm getting at, this relationship. Not trying to be completely useless here but whatever if this isn't interesting to you lol.

OpenStudy (anonymous):

hahah. it's not useless sir. still, thanks for the response. the more the reasons the better. @Kainui :)